OpenStudy (anonymous):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1) A:0.0392 B:0.0254 C:0.045

OpenStudy (abb0t):

To solve for heterozygous, you know that p = dominant q = recessive so 2pq, must be heterozygous. You already know, given by the information, that you have recessive, so q\(^2\) = \(\frac{1}{2500}\), solve for q, when p + q = 1 q = \(\sqrt{\frac{1}{2500}}\) p = 1-q p = ? But you don't know your entire population. So you will just leave it as "x"

OpenStudy (abb0t):

You want 2pq. Remember.

OpenStudy (lncognlto):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."